40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Jan 9 09:47:03 2010, Alberto, Update, LSC, Problems trying to lock the arms 
    Reply  Sat Jan 9 11:07:30 2010, Alberto, Update, LSC, Problems trying to lock the arms 
Message ID: 2492     Entry time: Sat Jan 9 11:07:30 2010     In reply to: 2491
Author: Alberto 
Type: Update 
Category: LSC 
Subject: Problems trying to lock the arms 


This morning I've been having problems in trying to lock the X arm.

The X arm's filter FM6 in the LSC screen starts blinking as it was halfloaded. Then the transmitted power drops from 1 to ~0.5 and eventually the arm loses lock.
To me it looked like a computer related issue. So I decide to reboot C1ISCEX by powercycling it.

That doesn't seem to have solved the problem. The X arm can get locked but TRX slowly moves between 0.2 and 1.

The X arm is now locked with TRX stable at ~1.

I think earlier on today I was having problems with running the alignment scripts from op540. Now I'm controlling the IFO from Rosalba and I can easily and stably lock all degrees of freedom.

I needed the X arm to be locked to align the auxiliary beam of the AbsL experiment to the IFO. To further stabilize TRX I increased the loop gain from 1 to 1.5.

Now the auxiliary beam is well aligned to the IFO and the beat is going through the PRC. I'm finally ready to scan the recycling cavity.

I also changed the gain of the PRC loop from -0.1 to -0.5.

ELOG V3.1.3-